The base of a triangular prism is an equilateral triangle with a perimeter of 42 cm. If the height of the prism is 10 cm, find the lateral area.
Answer by malakumar_kos@yahoo.com(315)


Base of a triangular prism = 42cms side of the triangle =42/3=14cms
height of the prism h = 10cms lateral surface area of the prism =circumference of the base.height= 2.22/7.14.10 =880 sq.cms